A turning fork vibrating at a frequency of 300 Hz is held over the top of a jar filled with water and fitted with a tap at the bottom. If the jar is 100 cm tall and the speed of sound of is 350 m/s, determine the possible resonance position (Neglect end correction)
1
Expert's answer
2020-11-10T06:59:35-0500
"\lambda=\frac{v}{f}=\frac{350}{300}=1.167\ m"

1st resonance at 


"d_1=\frac{116.7}{4}=29\ cm"

2nd resonance at 


"d_2=3\frac{116.7}{4}=88\ cm"